Problem 8 A large box of fuses contains 10% defectives. Four fuses are randomly selected from the box. Find: a) Probability that exactly one fuse is defective b) Probability that at least one fuse of the four selected is defective Now suppose these four sampled fuses are shipped to a customer before being tested. Assume the cost of fixing a shipment with defective fuses is given by C- 3Y2 where Y is the number of defectives in the shipment of four. Find the expected repair cost E(C)-E3Y
Answer #1

8)

Let Y denote the number of defective fuses. Then, Y~Bin(n=4,p=0.10)

P(У = y) = )0.1(1-0.1)4-у. У = 0.1, 2, 3, 4

So, E(Y)-np = 4(0.1) 0.4. Var(Y) = np(1-p) 4(0.1)(1-0.1)= 0.36

a)

Required probability = P(Y=1) = 4-1 10.11-0.10.2916

b)

Required probability = /4) 0.10(1-0, 1)4-0-0.3439

Now, 0.1728